OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

Second answer. You can get this by plugging in the points for x and y and seeing which ones get correct answers. Remember you have to test both points.

OpenStudy (anonymous):

Ok thanks everyone <3

OpenStudy (anonymous):

y-8=-3(x-5) -3x-15 +8=-7

OpenStudy (anonymous):

Use the slope formula to find the slope of the coordinates \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] And then use the numbers from the first coordinate along with the slope on this formula \[y-y _{1}=m(x-x _{1)}\]

OpenStudy (anonymous):

Thsnk spphire i will use that for my next one

OpenStudy (anonymous):

anytime :)